If the unit selling price is $20, the unit variable cost is $12 and the fixed costs are $180,000, what is the break-even point in units?

a. 15,000

b. 9,000

c. 22,500

d. 5,625
Answer #1

Contribution margin=Sales-Variable cost

=20-12=$8 per unit

Hence breakeven=Fixed cost/contribution margin

=180,000/8

=22500 units